A combined cycle gas turbine/vapor power plant uses the turbine exhaust as the energy source for the boiler. Each power system uses a single turbine. The gas power system is modeled as an ideal air-standard Brayton cycle. The vapor power system is modeled as an ideal Rankine cycle. Given specific operating conditions determine the temperature and pressure at each state, the rate of heat transfer in the boiler, the power output of each turbine, and the overall efficiency.

--Given Values--

Brayton Compressor Inlet: T1 (K) = 322

Brayton Compressor Inlet: P1 (kPa) = 190

Brayton Heat Exchanger Exit: T5 (K) = 530

Brayton: rp = 9.7

Brayton: rc = 2.45

Brayton: AV (m3/s) = 2.2

Rankine Turbine Inlet: T7 (C) = 480

Rankine Turbine Inlet: P7 (MPa) = 12

Rankine Condenser Inlet: P8 (kPa) = 50

a) Determine the specific enthalpy (kJ/kg) at the compressor inlet.

b) Determine the specific enthalpy (kJ/kg) at the combustor inlet.

c) Determine the specific enthalpy (kJ/kg) at the gas turbine inlet.

d) Determine the temperature (K) at the gas turbine inlet.

e) Determine the pressure (kPa) at the gas turbine inlet.

f) Determine the specific enthalpy (kJ/kg) at the gas turbine exit.

g) Determine the specific enthalpy (kJ/kg) of the gas at the heat exchanger exit.

h) Determine the mass flow rate of the air (kg/s) in the gas-power system.

i) Determine the net-power output (kW) of the Brayton cycle.

j) Determine the rate of heat transfer (kW) to the combustor.

k) Determine the rate of heat transfer (kW) to the water in the boiler.

l) Determine the specific enthalpy at the turbine inlet (kJ/kg).

m) Determine the specific entropy (kJ/kg-K) at the turbine inlet.

n) Determine the specific enthalpy at the condenser inlet (kJ/kg).

o) Determine the specific enthalpy at the pump inlet (kJ/kg).

p) Determine the specific enthalpy at the pump exit (kJ/kg).

q) Determine the mass flow rate of the water (kg/s) in the vapor-power system.

r) Determine the net power (kW) of the Rankine cycle.

s) Determine the thermal efficiency (%) of the combined power plant.
Answer #1

Ts=530k >Wy Heat exchanger I Brayton S cycle 4 2 190 kPa Wc A tares T1=322K 1 T1 7 To=4806 Claud [ Ranking Rankinę Cycle 12 Mspecific enthply at the gas 2 turbine inlet rc Tz Ts Tz = 2045× 530 }=1298.5k 3 Cpx 1298.5 h =1.005 X 1298.5 = 13ou.gg KJ kJ/3 heat exchanger Specific enthalpy of the exit at gas 2.4 To = oy Tix opt 322 x 9.74 T2= 616.29K hza Cpx Tz =1005X616.29 619.rate of heat transfer to water = mx Chu-hu = 2.64X ( 681.82 – 323.61) 45.67 KW Am At p= el 12Mpa Tau&oc specific enthply at thq = 340.54 +0.819 x(2645-2- 340.54) Tho 2228.05 Ks/vg Aus Specific enthply at pump inlet Ing=ht At p=50kPa [from steam team© Wivet Net power (kw) of the Rankine Cycle (WT-Wpxm I(haha)-(as hy) Jxmi [(32953 95:3 -2228005) - (12:30)] X0:321 - s IWnetA thumbs up, pls !!